• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sulkosky

    Log entry time 19:19:44 on July14,2003

    Entry number 103834

    Followups:

    keyword=Target Encoder Positions

    Figure 1 shows the updated target encoder positions for the second run period. The positions are similar to those for the first run period. The FSD switches have been checked and all have cleared expect for the Single "Long" 12C position. After the sweep of Hall A is complete, I will make an access to correct the position of the FSD switch.


    FIGURE 1